OpenStudy (sro03):

Hi I have this function of two variables: f(x,y)=x^2-y^2+xy And I need to find the gradient at the point (1,2). I tried an get the gradient at the point to (0,-11). It is true? Thanks in advance!

OpenStudy (kainui):

I'm getting something different, show your work and I can help

OpenStudy (sro03):

\[f(x,y)=x^2-y^2+xy\] P=(1,2) To find the gradient I first finde the derivate with respect to x and after with respect to y: \[gradient f(x,y)=f_x(x,y),f_y(x,y)\] <=> \[gradient f(x,y)=f_x(2x-2y+2), f_y(-3y^2+x)\] Now i insert the point P(1,2) : \[gradient f(2*1-2*2+2), (-3*2^2+1)\] = (0,-11) That's how I understand it :/

OpenStudy (kainui):

I see, so I think I see the issue here, when you compute \(f_x\) and \(f_y\) these are partial derivatives of \(f\) with respect to x and y, so when you write your gradient: \[\nabla f = \langle f_x, f_y \rangle = \left\langle \frac{\partial f}{\partial x}, \frac{\partial f}{\partial y} \right\rangle\] I think you're having some issues with calculating your partial derivatives, so I'll help walk you through it :D I'll do the partial with respect to x, then you try to do it with respect to y on your own and I'll help. \[\frac{\partial}{\partial x} ( x^2-y^2+xy) = \frac{\partial}{\partial x} ( x^2) +\frac{\partial}{\partial x} ( -y^2) +\frac{\partial}{\partial x} (xy) \] First step, I just look at each individual term sincethe derivative is a linear operator, you probably already know this, but it just lets us look at each term individually so we're not distracted. First term: \[\frac{\partial}{\partial x} ( x^2)=2x \] Second term: \[\frac{\partial}{\partial x} ( -y^2) =0\] Since y is independent of x, it's a constant with respect to a change in x and it's 0 just like any other constant would be. Last term: \[\frac{\partial}{\partial x} (xy)=y \] Similar situation here, y is a constant, but x isn't, so just like the derivative of \(\frac{\partial}{\partial x} (cx) = c\) that derivative is y. So: \[\frac{\partial f}{\partial x} = 2x+y\] So now try to calculate \(f_y\)

OpenStudy (sro03):

Then: \[\frac{ ∂f }{ ∂y }=x-2y\]

OpenStudy (kainui):

Perfect. Now calculate the gradient vector at the point and you're done

OpenStudy (sro03):

Do you get (4,-3) ? :/

OpenStudy (sro03):

\[2*1+2=4\] \[1-2*2=-3\] right?

OpenStudy (kainui):

Yup I got (4,-3) too

OpenStudy (sro03):

Thank you very much!! :)

OpenStudy (kainui):

Yeah you're welcome :)

OpenStudy (sro03):

I also have to find the \[D_uf\] I don't know the name for it in english but think you also call it \[D_uf\] ?

OpenStudy (sro03):

I get \[D_uf=\frac{ 11*\sqrt{5} }{5 }\] It is right?

OpenStudy (sro03):

Sorry you need some informations:

OpenStudy (sro03):

I have the direction R=(4,-2)

OpenStudy (sro03):

And the point (0,0) is a saddle piont, right?

OpenStudy (irishboy123):

you should compute \[\Large f_{xx}f_{yy}-f_{xy}f_{yx}\] \[\Large = f_{xx}f_{yy}-f^2_{xy}\] and switch x's and y's around... If it is \(<0\), it is a saddle point.